Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Joint Operation and Readiness Logbook, Contact: T.Haas, A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 546
Entry time: 16 Jul 2018, 14:18
|
Author: |
Adriano Violante |
Location: |
General |
Severity: |
Information |
Subject: |
Status report after dispatch meeting 13:00 |
|
|
SPB
- Mirrors installed. Lid closed. Las vacuum connection will be closed today. Then the electronics can be connected.
- Mock up pulse picker system ready
- Some crates in the tunnels will be switched on today after updates of the interlock definitions to be ready for photons.
- Train on demand technique is currently being discussed by FXE with DESY. SPB should be also involved.
SCS
- Some problems occur during the test of the PLCs, plus after firmware update new problems showed up. Patrick Gessler will organize a meeting with SCS and Bernard to improve the communication and identify the error sources.
SQS
- Tests ongoing. No relevant issues so far.
ITDM
- Working on new proposals format.
CAS
- XTD 6 open until the end of September.
- This week due to lack of resources by AE only the last slit and one PBLM will be commissioned. The rest will be done next week.
|